A large chunk of ice with a mass of 15.0 kg falls from a roof 8.00 m above the ground.
(a) What is the kinetic energy of the ice as it reaches the ground?
(b) What is its speed as it reaches the ground?
1
Expert's answer
2013-01-14T11:02:02-0500
The energy conservation law:
U + T = const
T - kinetic energy
U - potential energy
T = m*v^2/2
U = m*g*h
a) T = m*g*h = 15*8*10 =1200 J
b) v = Sqrt[2T/m]=Sqrt[2*1200/15] = 4*Sqrt[10] m/s = 12.6 m/s